# On tori periods of Weil representations of unitary groups

### v1

@@ 5.2. Global problem
By Theorem 5.4, the conditions in the theorem are necessary. Assuming these conditions, we will show the existence and uniqueness of a non-degenerate hermitian space $\mathrm{H}$ of dimension $n$ and class $\left[i\right]\in\Sigma_{\mathbf{E},\mathrm{H}}$, such that the period $\mathcal{P}_{i\left(L_{\mathbf{E}}^{1}\right),\alpha\circ i^{-1}}$ is not identically zero. Let us begin with uniqueness. By Theorem 2.8, for every $\mathrm{H}$, a choice of a class $\left[i\right]$ corresponds to an element $\lambda\in\mathbf{E}^{\times}/N_{L_{\mathbf{E}}/\mathbf{E}}\left(L_{\mathbf{E}}^{\times}\right)$, such that there exists an isomorphism of $\mathbf{K}$-hermitian spaces $r_{\lambda}\colon\mathrm{V}_{E,\lambda}\to\mathrm{H}$. If $\lambda_{1},\lambda_{2}\in\mathbf{E}^{\times}$ are such that there exist isomorphisms of hermitian spaces $r_{\lambda_{1}}\colon\mathrm{V}_{E,\lambda_{1}}\to\mathrm{H}$ and $r_{\lambda_{2}}\colon\mathrm{V}_{E,\lambda_{2}}\to\mathrm{H}$, such that the corresponding admissible embeddings $i_{\lambda_{1}}$ and $i_{\lambda_{2}}$ admit a non-zero period, then by condition (2) of Theorem 5.1, we must have that for every place $v$,
@@
and therefore for every place $v$, we have that $\lambda_{2}\lambda_{1}^{-1}\in N_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}/\mathbf{E}\otimes_{\mathbf{F}}\mathbf{F}_{v}}\left(\left(L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}\right)^{\times}\right)$. By the Hasse norm principle for quadratic extensions, this implies that $\lambda_{2}\lambda_{1}^{-1}\in N_{L_{\mathbf{E}}/\mathbf{E}}\left(L_{\mathbf{E}}^{\times}\right)$. Hence, $i_{\lambda_{1}}$ and $i_{\lambda_{2}}$ are conjugate. Suppose that $\mathrm{H}$ and $\mathrm{H}^{\prime}$ are both non-degenerate hermitian space of dimension $n$, and suppose that there exist $\lambda,\lambda^{\prime}\in\mathbf{E}^{\times}$ and isomorphisms of hermitian spaces $r_{\lambda}\colon\mathrm{V}_{\mathbf{E},\lambda}\to\mathrm{H}$ and $r_{\lambda^{\prime}}\colon\mathrm{V}_{\mathbf{E},\lambda^{\prime}}\to\mathrm{H}^{\prime}$, such that the corresponding admissible embeddings $i_{\lambda}$ and $i_{\lambda^{\prime}}$ admit a non-zero period. Then, as above, we get that $\lambda^{\prime}\lambda^{-1}\in N_{L_{\mathbf{E}}/\mathbf{E}}\left(L_{\mathbf{E}}^{\times}\right)$. This implies that $\mathrm{V}_{\mathbf{E},\lambda}$ is isomorphic to $\mathrm{V}_{\mathbf{E},\lambda^{\prime}}$, and therefore $\mathrm{H}$ is isomorphic to $\mathrm{H}^{\prime}$. We move to show the existence of a non-degenerate hermitian space $\mathrm{H}$ of dimension $n$ and a class $\left[i\right]\in\Sigma_{\mathbf{E},\mathrm{H}}$ with non-zero period $\mathcal{P}_{i\left(L_{\mathbf{E}}^{1}\right),\alpha\circ i^{-1}}$. We need to find an element $\lambda\in\mathbf{E}^{\times}$ such that for every place $v$, condition (2) of Theorem 5.4 holds. By writing $\mathbf{E}=\prod_{j=1}^{m}\mathbf{F}_{j}$, where $\mathbf{F}_{j}/\mathbf{F}$ is a finite field extension, the problem is reduced to the case where $\mathbf{E}/\mathbf{F}$ is a field extension, and that $L_{\mathbf{E}}/\mathbf{E}$ is a quadratic field extension. Consider the diagonal map $\Delta_{\mathbf{E}}\colon\mathbf{E}^{\times}/N_{L_{\mathbf{E}}/\mathbf{E}}\left(L_{\mathbf{E}}^{\times}\right)\to\left(\mathbf{E}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}\right)^{\times}/N_{L_{\mathbf{E}}/\mathbf{E}}\left(\left(L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}\right)^{\times}\right)$. This map has co-kernel
@@
Since the central $L$-function value $\mathcal{L}\left(\chi_{\mathrm{W}}^{-1}\circ N_{L_{\mathbf{E}}/\mathbf{K}}\cdot\alpha_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}}\right)$ is non-zero, we must have that the global root number $\prod_{v}\varepsilon_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}/\mathbf{E}\otimes_{\mathbf{F}}\mathbf{F}_{v}}\left(\alpha_{v,L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}}\cdot\chi^{-1}_{\mathrm{W},v}\circ N_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}/\mathbf{K}_{v}},\psi_{v},\delta\right)$ is $1$. Hence, the sequence $\left(\varepsilon_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}/\mathbf{E}\otimes_{\mathbf{F}}\mathbf{F}_{v}}\left(\alpha_{v,L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}}\cdot\chi^{-1}_{\mathrm{W},v}\circ N_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}/\mathbf{K}_{v}},\psi_{v},\delta\right)\right)_{v}$ lies in the image of $\Delta_{\mathbf{E}}$, and we can find $\lambda\in\mathbf{E}^{\times}$ as desired. ∎
@@ Appendix A Morphisms of norm one tori
In this appendix, we prove some technical statements regarding extensions of morphisms of $L_{E}^{1}$ to $L_{E}$. Let $F$ be an infinite field with characteristic different than $2$, and let $K/F$ be a quadratic étale algebra, equipped with an involution $x\mapsto x^{c}$ whose set of fixed points is $F$. Let $0\neq\delta\in K$ be a trace zero element. Any element $x\in K$ can be written in the form $x=a+b\delta$, where $a,b\in F$. We have that $\delta^{c}=-\delta$, and therefore $N_{K/F}\left(a+b\delta\right)=a^{2}-b^{2}\delta^{2}\in F$. Consider the map $j_{K/F}\colon K^{\times}\to K^{1}$ given by
@@
Then a simple computation yields $q_{F}\left(b\right)=b\delta$. Let $\mathrm{V}$ be an $n$-dimensional non-degenerate hermitian space over $K$. We are ready to prove our results.
@@
The equality (14) now follows from the fact that $j_{E}\left(1\pm b\delta\right)$ lies in $L_{E}^{1}$, and from the assumption that (12) holds for elements in $L_{E}^{1}$. If $m>1$, then by restricting to $L_{F_{j}}$, we get from the proof above that (12) holds for every $x\in L_{F_{j}}$, for every $1\leq j\leq m$. Using linearity, this implies that (12) holds for every $x\in L_{E}$. ∎
@@
Uniqueness follows from writing $\mathrm{m}_{\tau\left(x\right)}\otimes_{F}\operatorname{id}_{\bar{F}}=T^{-1}\circ\left(\mathrm{m}_{x}\otimes_{F}\operatorname{id}_{\bar{F}}\right)\circ T$ and applying both sides to $1\otimes 1\in L_{E^{\prime}}\otimes_{F}\bar{F}$. To show existence, first write $E=\prod_{j=1}^{m}F_{j}$, where $F_{j}/F$ is a field extension. Define for $b\in F_{j}$ with $\delta^{2}b^{2}\neq 1$,

### v2

@@ 5.2. Global problem
By Theorem 5.6, the conditions in the theorem are necessary. Assuming these conditions, we will show the existence and uniqueness of a non-degenerate hermitian space $\mathrm{H}$ of dimension $n$ and class $\left[i\right]\in\Sigma_{\mathbf{E},\mathrm{H}}$, such that the period $\mathcal{P}_{i\left(L_{\mathbf{E}}^{1}\right),\alpha\circ i^{-1}}$ is not identically zero.
Let us begin with uniqueness. By Theorem 2.8, for every $\mathrm{H}$, a choice of a class $\left[i\right]$ corresponds to an element $\lambda\in\mathbf{E}^{\times}/N_{L_{\mathbf{E}}/\mathbf{E}}\left(L_{\mathbf{E}}^{\times}\right)$, such that there exists an isomorphism of $\mathbf{K}$-hermitian spaces $r_{\lambda}\colon\mathrm{V}_{E,\lambda}\to\mathrm{H}$. If $\lambda_{1},\lambda_{2}\in\mathbf{E}^{\times}$ are such that there exist isomorphisms of hermitian spaces $r_{\lambda_{1}}\colon\mathrm{V}_{E,\lambda_{1}}\to\mathrm{H}$ and $r_{\lambda_{2}}\colon\mathrm{V}_{E,\lambda_{2}}\to\mathrm{H}$, such that the corresponding admissible embeddings $i_{\lambda_{1}}$ and $i_{\lambda_{2}}$ admit a non-zero period, then by condition (2) of Theorem 5.1, we must have that for every place $v$,
@@
and therefore for every place $v$, we have that $\lambda_{2}\lambda_{1}^{-1}\in N_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}/\mathbf{E}\otimes_{\mathbf{F}}\mathbf{F}_{v}}\left(\left(L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}\right)^{\times}\right)$. By the Hasse norm principle for quadratic extensions, this implies that $\lambda_{2}\lambda_{1}^{-1}\in N_{L_{\mathbf{E}}/\mathbf{E}}\left(L_{\mathbf{E}}^{\times}\right)$. Hence, $i_{\lambda_{1}}$ and $i_{\lambda_{2}}$ are conjugate.
Suppose that $\mathrm{H}$ and $\mathrm{H}^{\prime}$ are both non-degenerate hermitian space of dimension $n$, and suppose that there exist $\lambda,\lambda^{\prime}\in\mathbf{E}^{\times}$ and isomorphisms of hermitian spaces $r_{\lambda}\colon\mathrm{V}_{\mathbf{E},\lambda}\to\mathrm{H}$ and $r_{\lambda^{\prime}}\colon\mathrm{V}_{\mathbf{E},\lambda^{\prime}}\to\mathrm{H}^{\prime}$, such that the corresponding admissible embeddings $i_{\lambda}$ and $i_{\lambda^{\prime}}$ admit a non-zero period. Then, as above, we get that $\lambda^{\prime}\lambda^{-1}\in N_{L_{\mathbf{E}}/\mathbf{E}}\left(L_{\mathbf{E}}^{\times}\right)$. This implies that $\mathrm{V}_{\mathbf{E},\lambda}$ is isomorphic to $\mathrm{V}_{\mathbf{E},\lambda^{\prime}}$, and therefore $\mathrm{H}$ is isomorphic to $\mathrm{H}^{\prime}$.
We move to show the existence of a non-degenerate hermitian space $\mathrm{H}$ of dimension $n$ and a class $\left[i\right]\in\Sigma_{\mathbf{E},\mathrm{H}}$ with non-zero period $\mathcal{P}_{i\left(L_{\mathbf{E}}^{1}\right),\alpha\circ i^{-1}}$. We need to find an element $\lambda\in\mathbf{E}^{\times}$ such that for every place $v$, condition (2) of Theorem 5.6 holds. By writing $\mathbf{E}=\prod_{j=1}^{m}\mathbf{F}_{j}$, where $\mathbf{F}_{j}/\mathbf{F}$ is a finite field extension, the problem is reduced to the case where $\mathbf{E}/\mathbf{F}$ is a field extension, and that $L_{\mathbf{E}}/\mathbf{E}$ is a quadratic field extension. Consider the diagonal map $\Delta_{\mathbf{E}}\colon\mathbf{E}^{\times}/N_{L_{\mathbf{E}}/\mathbf{E}}\left(L_{\mathbf{E}}^{\times}\right)\to\left(\mathbf{E}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}\right)^{\times}/N_{L_{\mathbf{E}}/\mathbf{E}}\left(\left(L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}\right)^{\times}\right)$. This map has co-kernel
@@
Since the central $L$-function value $\mathcal{L}\left(\chi_{\mathrm{W}}^{-1}\circ N_{L_{\mathbf{E}}/\mathbf{K}}\cdot\alpha_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}}\right)$ is non-zero, we must have that the global root number $\prod_{v}\varepsilon_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}/\mathbf{E}\otimes_{\mathbf{F}}\mathbf{F}_{v}}\left(\alpha_{v,L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}}\cdot\chi^{-1}_{\mathrm{W},v}\circ N_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}/\mathbf{K}_{v}},\psi_{v},\delta\right)$ is $1$. Hence, the sequence
$\left(\varepsilon_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}/\mathbf{E}\otimes_{\mathbf{F}}\mathbf{F}_{v}}\left(\alpha_{v,L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}}\cdot\chi^{-1}_{\mathrm{W},v}\circ N_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}/\mathbf{K}_{v}},\psi_{v},\delta\right)\right)_{v}$
lies in the image of $\Delta_{\mathbf{E}}$, and we can find $\lambda\in\mathbf{E}^{\times}$ as desired.
Remark 5.9.
Similarly to Remark 5.2, by substituting $\mathrm{W}=\mathrm{W}_{\mathbf{F},1}^{\delta}$ and $\left(\chi_{\mathrm{W}},\chi_{\mathrm{V}}\right)=\left(\mu,\mu^{n}\right)$ as in Section 4.1.2, we get Theorem 1.5.
@@ Appendix A Morphisms of norm one tori
In this appendix, we prove some technical statements regarding extensions of morphisms of $L_{E}^{1}$ to $L_{E}$.
Let $F$ be an infinite field with characteristic different than $2$, and let $K/F$ be a quadratic étale algebra, equipped with an involution $x\mapsto x^{c}$ whose set of fixed points is $F$. Let $0\neq\delta\in K$ be a trace zero element. Any element $x\in K$ can be written in the form $x=a+b\delta$, where $a,b\in F$. We have that $\delta^{c}=-\delta$, and therefore $N_{K/F}\left(a+b\delta\right)=a^{2}-b^{2}\delta^{2}\in F$.
Consider the map $j_{K/F}\colon K^{\times}\to K^{1}$ given by
@@
Then a simple computation yields $q_{F}\left(b\right)=b\delta$.
Let $\mathrm{V}$ be an $n$-dimensional non-degenerate hermitian space over $K$. We are ready to prove our results.
@@
The equality (18) now follows from the fact that $j_{E}\left(1\pm b\delta\right)$ lies in $L_{E}^{1}$, and from the assumption that (16) holds for elements in $L_{E}^{1}$.
If $m>1$, then by restricting to $L_{F_{j}}$, we get from the proof above that (16) holds for every $x\in L_{F_{j}}$, for every $1\leq j\leq m$. Using linearity, this implies that (16) holds for every $x\in L_{E}$. ∎
@@
Uniqueness follows from writing $\mathrm{m}_{\tau\left(x\right)}\otimes_{F}\operatorname{id}_{\bar{F}}=T^{-1}\circ\left(\mathrm{m}_{x}\otimes_{F}\operatorname{id}_{\bar{F}}\right)\circ T$ and applying both sides to $1\otimes 1\in L_{E^{\prime}}\otimes_{F}\bar{F}$.
To show existence, first write $E=\prod_{j=1}^{m}F_{j}$, where $F_{j}/F$ is a field extension. Define for $b\in F_{j}$ with $\delta^{2}b^{2}\neq 1$,